OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 8, -14, 3+9i really need some help!

OpenStudy (anonymous):

@TuringTest could you help me out?

OpenStudy (anonymous):

@amistre64 could you help me out?

OpenStudy (anonymous):

@undeadknight26 could you help me out?

OpenStudy (mathmath333):

\(\large\tt \color{black}{x=8, x=-14,x= 3+9i}\) \(\large\tt \color{black}{therefore~~(x-8)(x+14)[x-(3+9i)]=0}\)

OpenStudy (mathmath333):

expand it u will get the equation

OpenStudy (anonymous):

alright, I mess up while working that out

OpenStudy (anonymous):

ill show you my work

OpenStudy (mathmath333):

bring it on

OpenStudy (anonymous):

(x-8)(x+14)(((x-3)-9i)((x-3)+9i) (x-8)(x+14)(x^2-6x-72) (x-8)(x^3+8x^2-156x+1008) x^4-220x^2+2256-8064 [final answer] This isn't one of the listed answers, so im thinking i went wrong somewhere...

OpenStudy (anonymous):

@undeadknight26 care to join?

undeadknight26 (undeadknight26):

Not good at this stuff..just learned it..

OpenStudy (anonymous):

listed answers A.)f(x)=x^4-11x^3+72x^2-606x+10800 B.)f(x)=x^4-303x^2+1212x-1080 C.)f(x)=x^4-11x^3-72x^2+606x-10800 D.)f(x)=x^4-58x^2+1212x-10800

OpenStudy (anonymous):

I just learned it too....

undeadknight26 (undeadknight26):

I...I have no idea im sorry.

OpenStudy (mathmath333):

calculation of (x-3)-9i)((x-3)+9i) is wrong

OpenStudy (mathmath333):

apply formula (a-b)(a+b)=a^2-b^2

OpenStudy (anonymous):

thank anyways undead and could you show mathmath?

OpenStudy (mathmath333):

(x-3)^2-(9i)^2=

OpenStudy (mathmath333):

do u know the value of i?

OpenStudy (anonymous):

yes imaginary numbers

OpenStudy (mathmath333):

so go on calculate

OpenStudy (anonymous):

so (x^2-9)+9?

OpenStudy (anonymous):

or would be just x^2?

OpenStudy (mathmath333):

x^2-6x+81+9

OpenStudy (anonymous):

oh ok i see what you did

OpenStudy (anonymous):

so would i take that and multiply it by (x-8)(x+14)(x^2-6x+81+9)

OpenStudy (mathmath333):

yes absolutely

OpenStudy (mathmath333):

x^2-6x+81+9=x^2-6x+90

OpenStudy (anonymous):

Alright let me work it and could you check it?

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

I got x^4-16x^3+70x^2+1196x-10080

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

alright i got it, you have to multiply (x-8) first the final answer was x^4-58x^2+1212x-10080

OpenStudy (anonymous):

thank you for taking the time and helping @mathmath333 I really appreciate it!

OpenStudy (mathmath333):

yes correct!!

OpenStudy (mathmath333):

welcome \(\Huge\tt \color{black}{\ddot\smile}\)